# VT-Former: A Transformer-based Vehicle Trajectory Prediction Approach For Intelligent Highway Transportation Systems

### v2

@@ IV-B1 Average Displacement Error (ADE)
IV-B2 Final Displacement Error (FDE)
quantifies the Euclidean distance between the ground truth final coordinates ($x^{T_{PH}},y^{T_{PH}}$) and the predicted final coordinates ($x^{\prime T_{PH}},y^{\prime T_{PH}}$) averaged over all available $N$ vehicles:
@@ IV-B2 Final Displacement Error (FDE)
IV-B3 Root Mean Squared Error (RMSE)
measures the square root of the average of the squared differences between the ground truth trajectory and the predicted trajectory at time step $t$ over all available $N$ vehicles:
@@ IV-B3 Root Mean Squared Error (RMSE)
IV-C Training Strategy and Hyperparameters
In accordance with the established conventions of prior research, our study maintains an observation horizon spanning 15 timesteps (equivalent to 3 seconds) and a prediction horizon extending to 25 timesteps (equivalent to 5 seconds) for all the datasets under consideration. Given the disparate recording rates of these datasets, a standardization process is undertaken to achieve comparability with previous research, necessitating the down-sampling of both datasets to a common frame rate of 5 frames per second.
Furthermore, aligned with established precedents, any instances where the complete trajectory data of a vehicle within a specified time window is unavailable are excluded from the datasets. This practice serves as a filtering and denoising mechanism. It is essential to acknowledge that, as discussed in Sec. IV-A, due to the relatively noisier nature of annotations within the CHD High-level and CHD Eye-level datasets in comparison to the NGSIM dataset, a notably higher volume of data points are omitted in the former datasets as part of this data refinement process.
@@ IV-C Training Strategy and Hyperparameters
V Results
TABLE I: Comprasion with SotA methods on NGSIM [12, 11]. The results are in meters.
RMSE
@@ V Results
CS-LSTM [14] 2.29 3.34 0.61 1.27 2.09 3.10 4.37
GRIP++ [34] 2.01 3.25 0.38 0.89 1.45 2.14 2.94
STA-LSTM [37] 1.89 3.16 0.37 0.98 1.71 2.63 3.78
DeepTrack [27] 2.01 3.21 0.47 1.08 1.83 2.75 3.89
Pisghu [4] 0.88 1.96 0.15 0.46 0.82 1.25 1.74
VT-Former 0.82 1.86 0.13 0.42 0.77 1.17 1.64
TABLE II: Comparison with SotA methods on CHD [28] High-angle dataset. The results are in pixels.
@@
VT-Former 25.95 87.21 7.60 17.35 22.90 27.97 66.39
TABLE III: Comparison with SotA methods on CHD [28] Eye-level dataset. The results are in pixels.
@@
VT-Former 34.88 100.59 6.71 17.24 28.70 45.86 82.00
V-A Comparison with State-of-the-Art Methods
In this section, we conduct a comparative analysis between VT-Former and other trajectory prediction methods, leveraging three extensive highway datasets discussed in Sec. IV-A: NGSIM [12, 11], CHD High-angle [28] and CHD Eye-level.
The results, presented in Tab. I, highlight performance metrics ADE, FDE, and RMSE on the NGSIM dataset. Values reported for GRIP++ [34] and STA-LSTM [37] are reported from [27]. VT-Former sets a new SotA on the NGSIM dataset, showcasing improvements of $6.8\%$ and $5.1\%$ in terms of ADE and FDE, respectively, surpassing the previous best-performing methods. In terms of RMSE, our model consistently demonstrates reduced errors, affirming its precision in trajectory prediction and establishing them as the leading choice for accurate trajectory prediction. The NGSIM dataset, comprising over 100,000 frames of data, offers a substantial volume of information, allowing the multi-head attention mechanism in our Transformer Predictor Module to be effectively trained. This, in turn, facilitates the model in capturing and learning intricate temporal patterns. Both CHD datasets are recently released, and as such, the majority of SotA models have not been evaluated on this specific dataset. To bridge this gap, we have used the available code repositories of leading SotA models, reconfiguring them for training and evaluation on the CHD datasets. As illustrated in Tab. II and Tab. III, our model ranks as the second-best performer in both the CHD High-angle and Eye-level datasets. This highlights its remarkable adaptability across a spectrum of environmental conditions and viewing angles, setting it apart from other models that tend to experience large performance degradation when transitioning between different settings. These outcomes underscore the synergistic effect of Graph Attentive Tokenization and the Transformer Predictor, which, when combined, capture complementary features, resulting in a versatile model suitable for deployment in various settings.
V-B Observation Horizon Analysis
@@
TABLE IV: VT-Former results using different observation horizons on NGSIM [12, 11] dataset. The results are in meters.
@@
$\mathbf{T_{OH}}$ ADE FDE 1s 2s 3s 4s 5s
1s 0.82 1.87 0.13 0.41 0.76 1.17 1.65
2s 0.88 1.97 0.15 0.45 0.82 1.26 1.75
3s 0.82 1.86 0.13 0.42 0.77 1.17 1.64
TABLE V: VT-Former results using different observation horizons on CHD High-angle dataset. The results are in pixels.
RMSE
$\mathbf{T_{OH}}$ ADE FDE 1s 2s 3s 4s 5s
1s 25.33 88.99 5.67 12.96 19.12 29.83 70.72
2s 25.90 87.90 6.40 14.62 20.56 29.44 70.05
3s 25.95 87.21 7.60 17.35 22.90 27.97 66.39
TABLE VI: VT-Former results using different observation horizons on CHD [28] Eye-level dataset. The results are in pixels.
RMSE
$\mathbf{T_{OH}}$ ADE FDE 1s 2s 3s 4s 5s
1s 21.86 66.28 5.42 12.69 18.81 26.67 53.05
2s 27.44 85.45 5.19 12.90 21.38 35.09 68.60
3s 34.88 100.59 6.71 17.24 28.70 45.86 82.00
In this section, we conduct a comprehensive investigation into the influence of different observation horizons to discern the optimal configuration for each dataset. The observation horizon is varied, ranging from 1 second to 3 seconds. Subsequently, we train and assess the performance of VT-Former on the NGSIM [12, 11], CHD High-angle [28], and CHD Eye-level [28] datasets. The results are presented in Tab. IV, Tab. III and Tab. V for the respective datasets. This specific assessment is especially important for applications that require faster prediction, such as collision avoidance or anomaly detection.
In the context of the NGSIM dataset, we have observed relatively minor variations in the accuracy of predicted trajectories when employing different observation horizons. These variations are at most $7.3\%$ and $5.3\%$ degradation in terms of ADE and FDE, respectively, as compared to the conventional 3-second observation setup. This particularly shows that even with a lower observation horizon, VT-Former is still able to accurately predict future trajectories. In the CHD High-angle dataset, a similar trend is observed, with the accuracy of predicted trajectories displaying consistency and minimal disparities among various observation horizons.
However, the results on the CHD Eye-level dataset, as depicted in Tab. VI, offer distinct insights. Here, it is evident that a shorter observation horizon is significantly more advantageous. Specifically, when deploying a 1-second observation horizon, we observe a remarkable enhancement of $37.3\%$ in FDE and $34.1\%$ in ADE compared to 3-second results. This outcome underscores the practicality of employing shorter observation horizons in the eye-level setup. This phenomenon can be attributed to the characteristics of the eye-level angle, wherein older observed frames are positioned farther from the camera. Consequently, a higher percentage of occlusions is encountered. Given that the annotations in the CHD datasets are extracted using deep learning algorithms, the increased occlusion contributes to a degradation in the quality of the extracted annotations, resulting in less accurate input data for trajectory prediction. Thus, a lower observation horizon translates to lower noise, and finally a more accurate prediction.
In summary, the empirical evidence presented in this study shows that, in practical, real-world high-angle and eye-level point-of-view scenarios marked by elevated occlusion rates among the detected vehicles, the adoption of shorter observation horizons is an effective strategy for decreasing the negative impact of noisy data. This approach not only leads to more accurate trajectory prediction outcomes but also offers the advantage of expeditious prediction. By reducing the temporal extent of observation, the model can effectively filter out the adverse effects of occlusion, which normally is seen in vehicles that are further from the camera, thus enhancing the precision of its predictions and enabling it to respond more swiftly. This finding underscores the practical utility of optimizing observation horizons in trajectory prediction models, particularly in applications where the timely and precise forecasting of object movements is important.
V-C Real-time Evaluation
Evaluating the real-time capabilities of VTP algorithms on embedded platforms is essential in ensuring the safety and reliability of these systems, as in many cases such as vehicle anomaly detection, they must make a split-second decision. By running trajectory prediction algorithms on an embedded board, developers can assess their performance in resource-constrained environments, and meet the low-latency requirements necessary to react promptly to changing conditions on the road. This process is not only important from the safety and efficiency perspectives but also helps bring cost-effective and energy-efficient technologies to a broader range of consumers.
(a) ROC curve with an AUC of 0.71
(b) PR curve with an AUC of 0.82
(c) FNR/FPR plots with EER of 0.32
Figure 3: Anomaly detection performance using VT-Former with 3-second input and 5-second prediction window.
VT-Former has been tested on Nvidia Jetson Orin AGX. The results can be seen in Tab. VII. As expected and discussed in Sec. IV-C it can be seen that the vehicles per frame of the NGSIM [11, 12] is higher than CHD Eye-level and CHD High-angle and the latency of NGSIM is significantly lower than each of the CHD datasets. As shown by the results, the latency of the model is 5.24, 69.89, and 48.09 milliseconds for NGSIM, CHD Eye-level, and CHD High-angle, respectively, indicating that the process can be done in real time. The results showcase that VT-Former can be adopted in the settings that require real-time processing, and where the speed of prediction and processing is of importance.
TABLE VII: VT-Former latency results on Jetson Orin AGX. The results are gathered using the test set of each dataset.
Dataset
Point
of View
Vehicles
per Frame
Latency
NGSIM [11, 12] Bird’s-eye 191 5.24
CHD [28] Eye-level 14 69.89
CHD [28] High-angle 21 48.09
VI VT-Former Usecase: Driving Anomaly Detection
Driving anomaly detection in an unsupervised setting refers to the process of identifying unusual or abnormal behavior or patterns in the operation of vehicles without the need for labeled training data. As anomaly detection is an open-set problem, with unsupervised setting, the model learns to predict the normal behavior in the training, and later in test and validation, when there is an anomalous behavior, the model’s prediction deviates from what is shown in the ground truth [3]. This type of anomaly detection is typically used for various purposes, such as ensuring the safety and security of vehicles, preventing accidents, and optimizing vehicle performance. The application of prediction-based unsupervised anomaly detection methodologies is a common practice in the field of pedestrian anomaly detection, particularly when focusing on pose-based anomaly detection [41, 3, 46]. In the context of driving anomaly detection, we extend these fundamental principles, as the trajectory of vehicles exhibits a close correlation with the pose exhibited by pedestrians.
In trajectory prediction, precision is imperative for constructing robust world models that anticipate naturalistic behaviors. Currently, VT-Former stands as the SotA on the NGSIM dataset [11, 12], exemplifying an accurate representation of these behaviors. The VT-Former model, trained on the NGSIM dataset, is capable of predicting normal driving behaviors. However, since it hasn’t been exposed to abnormal data during training it might struggle to accurately anticipate abnormal trajectories, resulting in significant deviations from ground truth. Hence, a higher ADE can serve as an effective metric for detecting anomaly scores.
While highways and roads are considered to be risky environments, there is a lack of proper anomalous driving datasets in this area. One of the main and valid reasons is that anomalies such as accidents are too dangerous for artificial creation. As such, normally an adversarial anomaly generator is utilized for assessment. We follow the same convention and build an anomaly dataset out of the NGSIM dataset [11, 12].
For assessment, we use the conventional metrics namely the Area Under The Receiver Operating Characteristic (AUC-ROC), the Area Under The Precision Recall Curve (AUC-PR), and the Equal Error Rate.
VI-A Adversarial Anomaly Trajectory Generation
Building upon discussions from prior sections, the absence of dedicated highway-based trajectory anomaly datasets presents a challenge in validating our anomaly detection methodologies. To circumvent this limitation, we adopt the recent advancements in adversarial anomaly generation as a benchmark for our proposed approach. Zhang et al., in their work [68], introduce an adversarial attack-based technique designed to craft realistic anomalous trajectories by perturbing standard trajectories within a dataset.
[68] uses six metrics to generate anomalies, such as ADE, FDE, deviation towards the left and right side of the lateral direction, and deviation towards the front and rear side of the longitudinal direction. The metrics for optimizing targeted disruptions are given as [24]:
$D(\alpha,R)=(p_{\alpha}-s_{\alpha})^{T}\cdot R(s_{\alpha+1},s_{\alpha}),$ (11)
where $\alpha$ represents the time frame, $R$ is a function that produces the unit vector towards a designated direction (lateral direction in our case), and $p$ and $s$ symbolize the predicted and actual vehicle positions, respectively.
Based on metrics, optimization has six different objective functions with several hard constraints based on the characteristics of the road and vehicle movements discussed in [68]. Finally, the optimization problem is designed to satisfy the objective and obey the constraints. To ensure adherence to the constraints, we adjust the perturbation, denoted ( $\Delta$), based on the historical trajectory ($H_{i}$) of the target vehicle $i^{th}$ and the constraint function ($C$). The optimization problem, as shown in the equation below, determines the maximum coefficient $\theta_{max}$ that scales the perturbation while satisfying all constraints for the vehicle of interest.
To evaluate our anomaly detection methodology, we produce anomalous data for the NGSIM dataset [12, 11] using the described adversarial trajectory generation technique.
VI-B Anomaly Detection Results
As discussed, the output of VT-Former is compared with the actual trajectory to classify each trajectory as normal or anomalous. As shown in Fig. 3, the results demonstrate an AUC-ROC of 0.71 and AUC-PR of 0.82 with an EER of 0.32. The results of the driving anomaly detector demonstrate moderately effective performance, with an AUC-ROC score of 0.71, indicating a reasonable ability to distinguish anomalies from non-anomalies. In contrast, the AUC-PR score of 0.82 reflects a strong balance between precision and recall in identifying anomalies. The EER at 0.32 signifies a specific threshold point where the detector equally balances false positives and false negatives. These findings suggest that the model is proficient in identifying anomalies and minimizing false positives, making it well-suited for applications where precision is crucial, such as driving anomaly detection in security or maintenance contexts. Fine-tuning the operating threshold may further optimize the model’s performance to align with specific application requirements.
VII Conclusion
@@
This work introduced VT-Former, an innovative trajectory prediction algorithm that leverages transformer architectures in conjunction with a novel tokenization scheme utilizing graph attention mechanisms denoted as Graph Attentive Tokenization (GAT). Extensive experimentation was conducted across three benchmark datasets illustrating SotA performance results, as well as assessing the model’s generalizability and robustness. Further experimentation aimed to demonstrate the real-time capabilities of VT-Former when deployed on an embedded platform, in addition to evaluating its performance under shorter observed horizons. Moreover, the study extended its contributions by implementing a driving anomaly detection approach built upon the foundation of VT-Former, thereby exemplifying a practical application of the proposed model.

### v4

@@ 4.2.1 Average Displacement Error (ADE)
4.2.2 Final Displacement Error (FDE)
FDE quantifies the Euclidean distance between the ground truth final coordinates ($x^{T_{PH}},y^{T_{PH}}$) and the predicted final coordinates ($x^{\prime T_{PH}},y^{\prime T_{PH}}$) averaged over all available $N$ vehicles:
@@ 4.2.2 Final Displacement Error (FDE)
4.2.3 Root Mean Squared Error (RMSE)
RMSE measures the square root of the average of the squared differences between the ground truth trajectory and the predicted trajectory at time step $t$ over all available $N$ vehicles:
@@ 4.2.3 Root Mean Squared Error (RMSE)
4.3 Training Strategy and Hyperparameters
In line with the conventions of previous research, our study utilizes an observation horizon of 15 timesteps and a prediction horizon of 25 timesteps (referred to as VT-FormerLH, where LH denotes Long History) across all datasets examined. Additionally, we define two more variations of VT-Former: the first version observes 10 timesteps and predicts 25 (VT-FormerMH, where MH stands for Medium History), and the second version observes 5 timesteps and predicts 25 (VT-FormerSH, where SH stands for Short History).
Given the disparate recording rates of these datasets, a standardization process is undertaken to achieve comparability with previous research, necessitating the down-sampling of both datasets to a common frame rate of 5 frames per second. For training the model for all datasets, Mean Squared Error (MSE) loss has been utilized.
@@ 4.3 Training Strategy and Hyperparameters
5 Results
Table 1: Comprasion with SotA methods (CS-LSTM [16], GRIP++ [46], STA-LSTM [49], DeepTrack [36], and Pishgu [3]) on NGSIM [13, 12]. VT-Former results are shown with different observation horizons of 15, 10, and 5 for VT-FormerLH, VT-FormerMH, and VT-FormerSH respectively. The results are in meters.
RMSE Params (k)
@@ 5 Results
CS-LSTM 2.29 3.34 0.61 1.27 2.09 3.10 4.37 191
GRIP++ 2.01 3.25 0.38 0.89 1.45 2.14 2.94 -
STA-LSTM 1.89 3.16 0.37 0.98 1.17 2.63 3.78 124
DeepTrack 2.01 3.21 0.47 1.08 1.83 2.75 3.89 109
VT-FormerLH 2.53 5.51 0.46 1.33 2.35 3.50 4.80 155
VT-FormerMH 2.24 5.15 0.37 1.13 2.10 3.26 4.62 141
VT-FormerSH 2.10 4.91 0.30 0.99 1.90 3.00 4.31 132
Table 2: Comparison with SotA methods on CHD [37] High-angle dataset. VT-Former results are shown with different observation horizons of 15, 10, and 5 for VT-FormerLH, VT-FormerMH, and VT-FormerSH respectively. The results are in pixels.
@@
VT-FormerLH 25.95 87.21 7.60 17.35 22.90 27.97 66.39
VT-FormerMH 25.90 87.90 6.40 14.62 20.56 29.44 70.05
VT-FormerSH 25.33 88.99 5.67 12.96 19.12 29.83 70.72
Table 3: Comparison with SotA methods on CHD [37] Eye-level dataset. VT-Former results are shown with different observation horizons of 15, 10, and 5 for VT-FormerLH, VT-FormerMH, and VT-FormerSH respectively. The results are in pixels.
@@
VT-FormerLH 34.88 100.59 6.71 17.24 28.70 45.86 82.00
VT-FormerMH 27.44 85.45 5.19 12.90 21.38 35.09 68.60
VT-FormerSH 21.86 66.28 5.42 12.69 18.81 26.67 53.05
5.1 Comparison with State-of-the-Art Methods
In this section, we conduct a comparative analysis between VT-Former and other trajectory prediction methods, leveraging three extensive highway datasets discussed in Sec. 4.1: NGSIM [13, 12], CHD High-angle [37] and CHD Eye-level [37]. Please note that VT-Former is only compared to deterministic models [3, 46, 36] and not models such as [58] that predict multiple future trajectories. Both CHD datasets were recently released, and as such, the majority of SotA models have not been evaluated on this specific dataset. To bridge this gap, we have used the available code repositories of leading SotA models, reconfiguring them for training and evaluation on the CHD datasets.
As discussed in Sec. 1, this exploratory study assesses the advantages and limitations of using transformers and graphs for SVTP. The results presented in Tab. 1, Tab. 2, and Tab. 3, along with the qualitative findings such as Fig. 4, collectively illustrate the potential of the VT-Former for highway trajectory prediction. Although the VT-Former did not achieve high results on the NGSIM dataset, it performed notably well as the second-best model on both the CHD High-angle and CHD Eye-level datasets. These findings indicate that while Graph Isomorphism Networks (GIN) and transformers are viable for SVTP, there is still potential for further refinement. More specifically, the data can lead to the hypothesis that the complexity of the graph significantly impacts the outcomes. Therefore, practices such as transitioning from fully connected graphs to more dynamic and optimized structures, along with implementing techniques such as neighborhood selection, could enhance performance.
Figure 4: Qualitative performance of VT-Former on CHD, featuring a scene with a merging lane from the left and a mild right curve. The angle between predicted, observed, and actual trajectories in samples A and B demonstrates VT-Former’s ability to predict the correct path with respect to road geometry. However, it struggles with the high acceleration of vehicle 2 in sample B. Samples are cropped for clarity.
5.2 Observation Horizon Analysis
In this section, we conduct a comprehensive investigation into the influence of different observation horizons to discern the optimal configuration for each dataset. The observation horizon is varied, including 5 timesteps for VT-FormerSH (translating to 1 second for 5Hz), 10 timesteps for VT-FormerMH (translating to 2 second for 5Hz) 15 timesteps for VT-FormerLH (translating to 3 second for 5Hz). Subsequently, we train and assess the performance of VT-Former on the NGSIM [13, 12], CHD High-angle [37], and CHD Eye-level [37] datasets. The results are presented in Tab. 1, Tab. 2, and Tab. 3 for the respective datasets. This specific assessment is especially important for applications that require faster prediction, such as collision avoidance or anomaly detection.
The results on all three datasets, as depicted in Tab. 1, Tab. 2, and Tab. 3, offer insights into the benefits of short-term observation horizon. Here, it is evident that a shorter observation horizon is more advantageous. For example, on the CHD Eye-level dataset, when deploying VT-FormerSH, we observe a remarkable enhancement of $37.3\%$ in FDE and $34.1\%$ in ADE compared to VT-FormerLH results. This outcome underscores the practicality of employing shorter observation horizons in the eye-level setup. Similar trends are seen for both NGSIM and CHD High-angle datasets. This phenomenon can be attributed to the nature of trajectory prediction. In longer trajectories, maintaining a high temporal resolution (detail at very small time scales) might be challenging due to computational constraints or model architecture limits. This loss of detail can lead to poorer predictions as the model might miss out on subtle cues that are more apparent in shorter, high-resolution trajectories. This particularly shows that with a lower observation horizon, VT-Former is able to predict future trajectories better.
In summary, the empirical evidence presented in this study shows that, in practical, real-world scenarios, the adoption of shorter observation horizons is an effective strategy for decreasing the negative impact of chaotic driving behavior. This approach not only leads to more accurate trajectory prediction outcomes but also offers the advantage of expeditious prediction. By reducing the temporal extent of observation, the model can effectively focus on details at smaller scales, maintaining a high temporal resolution, thus enhancing the precision of its predictions and enabling it to respond more swiftly. This finding underscores the practical utility of optimizing observation horizons in trajectory prediction models, particularly in applications where the timely and precise forecasting of object movements is important.
6 Ablation Study
In this section, we perform ablation testing to evaluate and compare the effect of higher data rate on the performance of the model, including the Observation Hozrion test same as Sec. 5.2. We utilize the NGSIM dataset as it offers a 10Hz sampling rate compared to conventional 5Hz sampling. The performance of VT-Former and Pishgu [3] with an observation horizon of 15 timesteps (translating to 1.5 seconds for 10Hz) and a prediction horizon of 25 timesteps (translating to 2.5 seconds for 10Hz) on the NGSIM dataset is presented in Tab. 4. VT-Former shows a $7.95\%$ improvement in ADE and $8.16\%$ improvement in FDE over Pishgu. A similar trend is observed in RMSE with the most pronounced improvement of $7.47\%$ from Pishgu’s 1.74 meters to 1.61 meters over a 2.5-second prediction horizon. This shows that at a higher data rate, VT-Former not only outperforms Pishgu but achieves higher accuracy with respect to RMSE at further timesteps.
@@ 6 Ablation Study
Table 4: Comparision of VT-Former and Pishgu on NGSIM dataset sampled at 10 Hz. VT-Former results are shown with different observation horizons of 15, 10, and 5 for VT-FormerLH, VT-FormerMH, and VT-FormerSH respectively. Pishgu results are from its updated repository. The results are in meters.
@@
Model ADE FDE 0.5s 1.0s 1.5s 2.0s 2.5s
Pisghu [3] 0.88 1.96 0.15 0.46 0.82 1.25 1.74
VT-FormerLH 0.81 1.80 0.14 0.42 0.76 1.16 1.61
VT-FormerMH 0.85 1.89 0.16 0.47 0.84 1.27 1.75
VT-FormerSH 0.79 1.78 0.13 0.41 0.76 1.16 1.62
Additionally, the impact of varying observation horizons, including 5 timesteps for VT-FormerSH (translating to 0.5 seconds for 10Hz), 10 timesteps for VT-FormerMH (translating to 1 second for 10Hz) 15 timesteps for VT-FormerLH (translating to 1.5 seconds for 10Hz), is assessed and detailed in Tab. 4. These results exhibit similar trends to those discussed in Sec. 5.2, where a shorter observation horizon correlates with improved performance in predicting future trajectories.
7 Conclusion
@@
This work introduced VT-Former, an innovative trajectory prediction algorithm that leverages transformer architectures in conjunction with a novel tokenization scheme utilizing graph attention mechanisms denoted as Graph Attentive Tokenization (GAT). Extensive experimentation was conducted across three benchmark datasets illustrating the performance of the model, as well as assessing the model’s generalizability and robustness. Further experimentation has been conducted to evaluate VT-Former’s performance under shorter observed horizons. This study underscores the potential of combining graphs with transformers, opening a new path for future research in the field.